OpenStudy (anonymous):

2x+y=-4 and 4x-2y=8 how many solutions?

OpenStudy (anonymous):

ans is x =0 , y=-4

OpenStudy (anonymous):

2x + y = -4 y = -2x - 4 4x - 2y = 8 2y = 4x - 8 y = 2x -4 1 solution

OpenStudy (anonymous):

so two solution

OpenStudy (anonymous):

It's one solution (0, -4), right?

OpenStudy (anonymous):

no its two bcas two variable so two solution ...got it ...

OpenStudy (anonymous):

No, it's one solution because (0,-4) is one order pair.

OpenStudy (anonymous):

how do i know its on ordered pair?

OpenStudy (anonymous):

Okay, so when you work out a solution to your problem and you end up with the sum of 2 variables, those are your ordered pair. For example, If a=2 b=3 Your order pair would be (2,3) Basically, you have (a,b) or whatever given variable Then you work out your equation, then plug the numbers into the variables. Get it?